# Context-Free Synthetic Data Mitigates Forgetting

### v1

@@ 1 Introduction
Figure 1: We finetune Olmo-1B Groeneveld et al. (2024) model on MetaMathQA Yu et al. (2023) dataset with the aim of improving GSM8K accuracy while maintaining it’s pre-existing (i.e., pretrained) abilities (kindly refer to Sec 3.1 for details). Our method CFS, augments the downstream data with context-free synthetic data (Sec 2) and performs better than the considered baselines. Pretrain-Aug augments MetaMathQA with pretraining data, LoRA trains a low-rank adaptation, $l_{2}$ regularization regularizes model towards it’s initialization and Wise-FT does post-hoc model averaging of Finetuned and Base.
Forgetting occurs because model weights shift during fine-tuning; attenuating this shift attenuates forgetting, while also possibly attenuating the new-task gains from fine-tuning. This is the realization underlying several existing methods to mitigate forgetting in the data-oblivious setting; these include adding an $\ell_{2}$ regularization penalty to the change in weights Kirkpatrick et al. (2017); Kumar et al. (2023), deliberately using LoRA while freezing weights in the main model Hu et al. (2022); Biderman et al. (2024), freezing subsets of parameters Chen et al. (2024); Panda et al. (2024), model-averaging Wortsman et al. (2021), selecting easy samples during fine-tuning Sanyal et al. (2025) etc.
@@
  • We first demonstrate the efficacy of this method when adding a new task to a pre-trained only model; specifically, we consider the Olmo-1B Groeneveld et al. (2024) model [^1] We chose the Olmo-1B model because it’s pretraining data is actually available, and we wanted to benchmark the efficacy of our data-oblivious method against classical “replay based” continual learning approaches Rolnick et al. (2019) that need and use pre-training data., which is then fine-tuned on the MetaMathQA dataset Yu et al. (2023) in an attempt to improve its performance on GSM8K. We show that our method works better than other benchmark data-oblivious methods like $\ell_{2}$ regularization, LoRA and model averaging, as well as data-aware methods like replay. See Figure 1 and Section 3.1.
  • We then turn our attention to reasoning models, and ask the question of whether fine-tuning degrades reasoning capacity (it does), and how to prevent this forgetting. Specifically, we investigate the effect on math reasoning performance of R1-Distill-Llama8B DeepSeek-AI (2025) when it is fine-tuned on the medical MedReason Wu et al. (2025) dataset. Here again we show that our method improves on other benchmark data-oblivious methods in mitigating forgetting. See Figure 2 and Section 3.2 for details.
@@
Figure 2: We finetune R1-Distill-Llama-8B DeepSeek-AI (2025) model on MedReason Wu et al. (2025) dataset with the aim of improving it’s medical abilities, while maintaing it’s reasoning performance (kindly refer to Sec 3.2 for details). Our method CFS, augments the downstream data with context-free synthetic data (Sec 2) and performs better than the considered baselines, namely LoRA which trains a low-rank adaptation, which regularizes model towards it’s initialization and Wise-FT which does post-hoc model averaging of Finetuned and Base.
@@ 1.1 Related work
Catastrophic forgetting  Wang et al. (2024); Zheng et al. (2025) has a rich literature, highlighting the importance of the problem. Methods to mitigate forgetting, termed as continual learning methods can broadly be classified as data-oblivious approaches (i.e., they don’t assume access to prior data used to train the model) and data-dependent approaches (assume access to some subset of the data). Our focus in this paper is on the data-oblivious setting and we review relevant works in this section.
@@
Regularization-based approaches A class of well known and intuitive approaches for mitigating forgetting in data-oblivious setting constrain the learned model weights to be close to the initial model weights in a suitable metric. A simple idea is to constraint the learned weights to be close in the $\ell_{2}$ norm Kumar et al. (2023); Kirkpatrick et al. (2017). LoRA Hu et al. (2022) enforces a low rank difference between the weight matrices of the learned weights and the initial weights.  Biderman et al. (2024) shows that LoRA indeed mitigates forgetting, though can also hurt effective adaptation to new tasks. Another line of work doesn’t constraint the problem while training, but instead post-hoc averages (or a smarter convex combination) the learned model weights with initial weights to tradeoff between learning and forgetting Lubana et al. (2021); Wortsman et al. (2021); Ilharco et al. (2023); Lin et al. (2023); Kleiman et al. (2025).
@@
Synthetic data-based approaches A line of work in data-dependent approaches focuses on caching samples from previously seen tasks and augmenting using these samples when finetuning for the new task de Masson D’Autume et al. (2019); Rolnick et al. (2019). Since the data-oblivious setting is a more realistic scenario, prior work propose using a generative model to stand-in for the previously seen data. Specifically, they jointly train both a generative model and a classifier with the generative model standing in as a proxy for previously seen tasks  Wu et al. (2018); Kemker and Kanan (2018); Smith et al. (2021); Yin et al. (2020).  Huang et al. (2024) proposes a similar setup using language model. It prompts the language model to synthesis examples similar to ones seen in the previous tasks.
@@
Other data-agnostic methods Apart from theses techniques,  Sanyal et al. (2025) explores reweighing data for mitigating forgetting. Specifically, they use the base model’s own likelihood for finetuning to re-weigh samples before finetuning. They up-weigh easy samples and show that this mitigates forgetting. Chen et al. (2024); Panda et al. (2024) also propose data-agnostic methods for continual learning which leverage gradient and other information to select a subset of parameters to update while finetuning. Similar to our work, Yang et al. mitigates forgetting by a KL divergence term between the output token probability of the base model and the learned model. We consider a more general setup where we minimize KL-divergence over the set of all strings.
@@
While recent works on aligning language models  Shao et al. (2024); DeepSeek-AI (2025) have extensively explored training on model’s own responses, with some verification in loop, how model generated synthetic data can help mitigate forgetting is still under-explored. We show that particularly context-free generations are helpful for mitigating forgetting.
@@ 2 Context-free synthetic data
Setup Let us denote, as is the convention, a language model by $p_{\theta}$, with $\theta$ being the weights; in particular $p_{\theta}(x)$ denotes the probability [^2] In particular, the standard notion of LLM probability of a sequence is the product of the logits the model assigns to each token in the sequence. the model assigns to a string $x$. With this notation, pretraining a model on a dataset $\mathcal{D}$ can be written as
$\min_{\theta}\quad E_{x\sim\mathcal{D}}\,\left[\,-\log p_{\theta}(x)\,\right]\quad\quad\text{starting from rand init}$ (1)
@@
$\min_{\theta}\quad E_{(x,y)\sim\mathcal{F}}\,\left[\,-\log p_{\theta}(y|x)\,\right]\quad+\quad\lambda\,\text{KL}(p_{\theta^{*}}\|p_{\theta})\quad\quad\text{starting from $\theta^{*}$}$ (2)
where KL$(p_{\theta^{*}}\|p_{\theta})$ stands for the Kullback-Liebler divergence between the original model $p_{\theta^{*}}$ and the new model $p_{\theta}$ (and $\lambda$ is a penalty parameter). Of course, the problem with this conceptual approach is that it is not clear what a term like KL$(p_{\theta^{*}}\|p_{\theta})$ operationally means, or how to calculate it.
@@
Note now the first term in the last equation above does not depend on $\theta$; thus, minimizing KL$(p_{\theta^{*}}\|p_{\theta})$ is equivalent to minimizing the second term. Putting this back into (2) yields the following
$\min_{\theta}\quad E_{(x,y)\sim\mathcal{F}}\,\left[\,-\log p_{\theta}(y|x)\,\right]\quad+\quad\lambda\,E_{x\sim p_{\theta^{*}}}\left[-\log p_{\theta}(x)\right]$
@@
(1) $\ell_{2}$ penalization: while our penalty term KL$(p_{\theta^{*}}\|p_{\theta})$ penalizes shift in distributions, one could instead directly penalize a shift in the parameters themselves, i.e. have a penalty term $\|\theta-\theta^{*}\|^{2}_{2}$. One disadvantage of this method is that it needs to keep two sets of model weights ($\theta$ and $\theta^{*}$) in memory while training, while ours does not; one disadvantage of our method is that synthetic data generation may be slow because it is sequential.
(2) Conditional generation: A natural synthetic data approach one may think of is to augment the fine-tuing data with condtional generation; that is, for every $(x,y)$ in $\mathcal{F}$, generate a $\widehat{y}\sim p_{\theta^{*}}(y|x)$ by giving the $x$ as an input context to the model $\theta^{*}$; and apply the standard fine-tuning loss on both the original $(x,y)$ and the new $(x,\widehat{y})$. We show that this performs quite poorly; this is because it is pushing the model in different directions for the exact same input context.
(3) Replay: The idea of experience replay Rolnick et al. (2019), in our context, involves retaining some portion of the past data used to train $\theta^{*}$, and adding it in during fine-tuning. Of course this is not data-oblivious, and is often not applicable in modern “open weights but not open data" regimes. For one the models in this paper - Olmo-1B - the pretraining data is available, and we compare against this for that model; the other two models are “weights only" models and hence we cannot implement replay for them.
@@
Model Input Prompt Examples for Context-Free Generation
Olmo-1B <|endoftext|> The coronavirus pandemic has caused changes in people’s life as never before, with many people avoiding public spaces and adhering to …
@@ 3 Experiments
Context-free synthetic data Recall that a key step in our method is context-free data generation. For context-free generations we prompt the model with it’s bos_token,i.e., beginning of sentence token. Operationally, for the two models we investigate in this paper, Table 1 shows the input prompts used for the generations. We use a sampling temperature of 1.0 unless otherwise specified, and top-p of 0.95 (we ablate on the temperature hyperparameter in our experiments). We use vLLM Kwon et al. (2023) for generation. Another choice for generating synthetic samples is contextual generations. For contextual generations, we generate model responses for each input prompt in the finetuning dataset and use the model’s responses (along with the input prompts) as the augmented samples. For contextual-generations we use sampling temperature of 0.6 and probability threshold (top-p) of 0.95. For pretrained-only models, we also explore augmenting with their pretraining corpus. Specifically, for Olmo we subsample 400K rows (equal to number of samples in MetaMathQA) from it’s pretraining corpus Dolma dataset Soldaini et al. (2024).
@@
Methods and models The tables in this paper study the performance of models developed using the following methods (all of which, like ours, are data-oblivious - except P):
  • Base
@@
    This refers to the zero-shot performance of the base model (Olmo-1B for pre-trained-only or R1-Distill-Llama-8B for reasoning)
  • FT
@@
    This refers to the model that results when standard supervised fine-tuning is applied to the Base model.
  • P
@@
    We chose the Olmo-1B model because we have access to its pretraining data, which allows us to compare against the replay approach Rolnick et al. (2019) – i.e. training a model on a combination of the standard fine-tuning loss on the finetuning dataset, and pretraining loss on a (random subset of) the pretraining data.
  • CS
@@
    This refers to contextual generation – i.e. for each sample in the fine-tuning dataset, make a new sample which contains the same input context but now paired with what the old model’s generated answer to that input. Fine-tune on all samples in this augmented dataset.
  • CFS
    This is our method, based on context-free synthetic data.
  • LoRA
@@
    This refers to the model that results when LoRA Hu et al. (2022) is used in the standard supervised fine-tuning stage; this is based on the paper Biderman et al. (2024) which shows that LoRA can mitigate forgetting; this is thus a baseline method we compare against.
  • Wise-FT
    Taken from Wortsman et al. (2021), this is another baseline method, based on model averaging. In particular, it advocates first doing standard fine-tuning, and then devising a weighted combination of $\alpha\times$ original model weights and $(1-\alpha)\times$ new model weights.
  • $\ell_{2}$
@@
    As described in Kumar et al. (2023); Kirkpatrick et al. (2017), this is our final baseline; it advocates adding the $\ell_{2}$ distance between the old model weights and the new model weights into the loss. Note that this involves needing to store two full models in memory during training, which becomes cumbersome for large models.
@@
We wish to investigate the effect of different augmentation sources on model performance. We use 1:1 mix of finetuning data and augmented samples, unless otherwise specified. For a fair comparison between different mix of datasets, we control for number of gradient steps, i.e. the size of dataset is inversely proportional to the number of epochs we use it for. We use the AdamW optimizer with a cosine learning rate schedule with a peak learning rate of $65e-$ 6 and warmup steps of $3\%$ of total training steps and train for 2-4 epochs (see Supp B). We have an effective batch size of 128 (following  Wu et al. (2025); Yuan et al. (2025)). See Supp B for additional training details.

### v2

@@ 1 Introduction
Figure 1: We finetune Olmo-1B (Groeneveld et al., 2024) model on MetaMathQA (Yu et al., 2023) dataset with the aim of improving GSM8K accuracy while maintaining it’s pre-existing (i.e., pretrained) abilities (kindly refer to Sec 3.1 for details). Our method CFS, augments the downstream data with context-free synthetic data (Sec 2) and performs better than the considered baselines. Pretrain-Aug augments MetaMathQA with pretraining data, LoRA trains a low-rank adaptation, $l_{2}$ regularization regularizes model towards it’s initialization and Wise-FT does post-hoc model averaging of Finetuned and Base.
Forgetting occurs because model weights shift during fine-tuning; attenuating this shift attenuates forgetting, while also possibly attenuating the new-task gains from fine-tuning. This is the realization underlying several existing methods to mitigate forgetting in the data-oblivious setting; these include adding an $\ell_{2}$ regularization penalty to the change in weights (Kirkpatrick et al., 2017; Kumar et al., 2023), deliberately using LoRA while freezing weights in the main model (Hu et al., 2022; Biderman et al., 2024), freezing subsets of parameters (Chen et al., 2024; Panda et al., 2024), model-averaging (Wortsman et al., 2021), selecting easy samples during fine-tuning (Sanyal et al., 2025) etc.
@@
Our work doesn’t claim novelty in proposing synthetic data augmentation for mitigating forgetting. Our aim with this work is to show that our simple context-free generations from a language model are (surprisingly) better suited for mitigating forgetting (in context of LLMs) than intuitive naive choices including : (a) domain-specific contextual generations : here the input context for generation are given by the input prompts of the finetuning data itself. (b) pretraining data : for settings where we consider pretrained-only model with open access to it’s pretraining data, our results show that context-free generations outperform augmentation with the pretraining data as well (which is equivalent to data-replay augmentation). We highlight that work is focused on evaluating the effectiveness of different sources of data for mitigating forgetting and doesn’t address practical challenges like cost of generation etc.
Figure 2: We finetune R1-Distill-Llama-8B (DeepSeek-AI, 2025) model on MedReason (Wu et al., 2025) dataset with the aim of improving it’s medical abilities, while maintaing it’s reasoning performance (kindly refer to Sec 3.2 for details). Our method CFS, augments the downstream data with context-free synthetic data (Sec 2) and performs better than the considered baselines, namely LoRA which trains a low-rank adaptation, which regularizes model towards it’s initialization and Wise-FT which does post-hoc model averaging of Finetuned and Base.
@@
  • We first demonstrate the efficacy of this method when adding a new task to a pre-trained only model; specifically, we consider the Olmo-1B (Groeneveld et al., 2024) model [^1] We chose the Olmo-1B model because it’s pretraining data is actually available, and we wanted to benchmark the efficacy of our data-oblivious method against classical “replay based” continual learning approaches (Rolnick et al., 2019) that need and use pre-training data., which is then fine-tuned on the MetaMathQA dataset (Yu et al., 2023) in an attempt to improve its performance on GSM8K. We show that our method works better than other benchmark data-oblivious methods like $\ell_{2}$ regularization, LoRA and model averaging, as well as data-aware methods like replay. See Figure 1 and Section 3.1.
  • We then turn our attention to reasoning models, and ask the question of whether fine-tuning degrades reasoning capacity (it does), and how to prevent this forgetting. Specifically, we investigate the effect on math reasoning performance of R1-Distill-Llama8B (DeepSeek-AI, 2025) when it is fine-tuned on the medical MedReason (Wu et al., 2025) dataset. Here again we show that our method improves on other benchmark data-oblivious methods in mitigating forgetting. See Figure 2 and Section 3.2 for details.
@@ 1.1 Related work
Catastrophic forgetting  (Wang et al., 2024; Zheng et al., 2025) has a rich literature, highlighting the importance of the problem. Methods to mitigate forgetting, termed as continual learning methods can broadly be classified as data-oblivious approaches (i.e., they don’t assume access to prior data used to train the model) and data-dependent approaches (assume access to some subset of the data). Our focus in this paper is on the data-oblivious setting and we review relevant works in this section.
@@
Regularization-based approaches A class of well known and intuitive approaches for mitigating forgetting in data-oblivious setting constrain the learned model weights to be close to the initial model weights in a suitable metric. A simple idea is to constraint the learned weights to be close in the $\ell_{2}$ norm (Kumar et al., 2023; Kirkpatrick et al., 2017). LoRA (Hu et al., 2022) enforces a low rank difference between the weight matrices of the learned weights and the initial weights.  (Biderman et al., 2024) shows that LoRA indeed mitigates forgetting, though can also hurt effective adaptation to new tasks. Another line of work doesn’t constraint the problem while training, but instead post-hoc averages (or a smarter convex combination) the learned model weights with initial weights to tradeoff between learning and forgetting (Lubana et al., 2021; Wortsman et al., 2021; Ilharco et al., 2023; Lin et al., 2023; Kleiman et al., 2025).
@@
Synthetic data-based approaches A line of work in data-dependent approaches focuses on caching samples from previously seen tasks and augmenting using these samples when finetuning for the new task (de Masson D’Autume et al., 2019; Rolnick et al., 2019). Since the data-oblivious setting is a more realistic scenario, prior work propose using a generative model to stand-in for the previously seen data. Specifically, they jointly train both a generative model and a classifier with the generative model standing in as a proxy for previously seen tasks  (Wu et al., 2018; Kemker and Kanan, 2018; Smith et al., 2021; Yin et al., 2020).  (Huang et al., 2024) proposes a similar setup using language model. It prompts the language model to synthesis examples similar to ones seen in the previous tasks.
@@
Other data-agnostic methods Apart from theses techniques,  (Sanyal et al., 2025) explores reweighing data for mitigating forgetting. Specifically, they use the base model’s own likelihood for finetuning to re-weigh samples before finetuning. They up-weigh easy samples and show that this mitigates forgetting. (Chen et al., 2024; Panda et al., 2024) also propose data-agnostic methods for continual learning which leverage gradient and other information to select a subset of parameters to update while finetuning. Similar to our work, (Yang et al., ) mitigates forgetting by a KL divergence term between the output token probability of the base model and the learned model. We consider a more general setup where we minimize KL-divergence over the set of all strings.
@@
While recent works on aligning language models  (Shao et al., 2024; DeepSeek-AI, 2025) have extensively explored training on model’s own responses, with some verification in loop, how model generated synthetic data can help mitigate forgetting is still under-explored. We show that particularly context-free generations are helpful for mitigating forgetting.
Model Input Prompt Examples for Context-Free Generation
Olmo-1B <|endoftext|> The coronavirus pandemic has caused changes in people’s life as never before, with many people avoiding public spaces and adhering …
[…]
@@ 2 Context-free synthetic data
Setup Let us denote, as is the convention, a language model by $p_{\theta}$, with $\theta$ being the weights; in particular $p_{\theta}(x)$ denotes the probability [^2] In particular, the standard notion of LLM probability of a sequence is the product of the probabilities the model assigns to each token in the sequence. the model assigns to a string $x$. With this notation, pretraining a model on a dataset $\mathcal{D}$ can be written as
$\min_{\theta}\quad E_{x\sim\mathcal{D}}\,\left[\,-\log p_{\theta}(x)\,\right]\quad\text{starting from rand init}$ (1)
@@
$\min_{\theta}\quad E_{(x,y)\sim\mathcal{F}}\,\left[\,-\log p_{\theta}(y|x)\,\right]\;+\;\lambda\,\text{KL}(p_{\theta^{*}}\|p_{\theta})$ (2)
[…]
@@
[…]
$\min_{\theta}\hskip 5.0ptE_{(x,y)\sim\mathcal{F}}\left[\,-\log p_{\theta}(y|x)\,\right]\;+\;\lambda\,E_{x\sim p_{\theta^{*}}}\left[-\log p_{\theta}(x)\right]$
@@
[…]
(3) Replay: The idea of experience replay (Rolnick et al., 2019), in our context, involves retaining some portion of the past data used to train $\theta^{*}$, and adding it in during fine-tuning. Of course this is not data-oblivious, and is often not applicable in modern “open weights but not open data" regimes. For one the models in this paper - Olmo-1B - the pretraining data is available, and we compare against this for that model; the other two models are “weights only" models and hence we cannot implement replay for them.
@@
Pre-existing Tasks
Commonsense MMLU BB Hard AGIEval Avg. GSM8K
Standard Base 50.35 24.36 25.56 17.73 29.50 1.59
FT 40.89 23.06 11.49 18.77 23.55 29.49
Data-augmentation CS 42.97 24.16 14.77 18.28 25.05 19.71
P 46.67 23.01 10.69 18.56 24.73 15.09
CFS 46.96 23.29 23.19 19.52 28.24 29.34
Weight-Space LoRA $(r=256)$ 48.41 23.71 21.87 18.72 28.18 20.17
$\ell_{2}$ regularization 47.43 23.03 18.92 19.76 27.28 28.43
@@
Table 2: This table studies the pretrained-only (i.e. Olmo-1B, MetaMathQA, GSM8K) setup described in Section 3.1. We see that standard finetuning FT is better than on Base the downsream task (GSM8K) but worse on the pre-existing tasks; this is forgetting. The next three rows report numbers with data-augmentation methods–including our method CFS. The table shows that CFS is much more effective at mitigating forgetting as compared to the other two data-augmentation-based methods P and CS. Next we compare our method (CFS) to the popular weight-space approaches LoRA, $\ell_{2}$ and Wise-FT. Each of the baselines have hyper-parameters which we optimized over (see appendix); we report their best numbers here. We similarly ablate over the amount of context-free samples in our method (see Table 3) and report the best numbers here. While all these methods mitigate forgetting, our CFS seems the most effective.
@@ 3 Experiments
Context-free synthetic data Recall that a key step in our method is context-free data generation. For context-free generations we prompt the model with it’s bos_token,i.e., beginning of sentence token. Operationally, for the two models we investigate in this paper, Table 1 shows the input prompts used for the generations. We use a sampling temperature of 1.0 unless otherwise specified, and top-p of 0.95 (we ablate on the temperature hyperparameter in our experiments). We use vLLM (Kwon et al., 2023) for generation. Another choice for generating synthetic samples is contextual generations. For contextual generations, we generate model responses for each input prompt in the finetuning dataset and use the model’s responses (along with the input prompts) as the augmented samples. For contextual-generations we use sampling temperature of 0.6 and probability threshold (top-p) of 0.95. For pretrained-only models, we also explore augmenting with their pretraining corpus. Specifically, for Olmo we subsample 400K rows (equal to number of samples in MetaMathQA) from it’s pretraining corpus Dolma dataset (Soldaini et al., 2024).
@@
Methods and models The tables in this paper study the performance of models developed using the following methods (all of which, like ours, are data-oblivious - except P, which in our setup is equivalent to data-replay):
@@
Base: This refers to the zero-shot performance of the base model (Olmo-1B for pre-trained-only or R1-Distill-Llama-8B for reasoning)
@@
FT: This refers to the model that results when standard supervised fine-tuning is applied to the Base model.
@@
P: We chose the Olmo-1B model because we have access to its pretraining data, which allows us to compare against the data replay approach (Rolnick et al., 2019) – i.e. training a model on a combination of the standard fine-tuning loss on the finetuning dataset, and pretraining loss on a (random subset of) the pretraining data. Since Olmo-1B is a pretrained-only model, this approach is exactly equivalent to data-replay.
@@
CS: This refers to contextual generation – i.e. for each sample in the fine-tuning dataset, make a new sample which contains the same input context but now paired with what the old model’s generated answer to that input. Fine-tune on all samples in this augmented dataset.
CFS: This is our method, based on context-free synthetic data.
@@
LoRA: This refers to the model that results when LoRA (Hu et al., 2022) is used in the standard supervised fine-tuning stage; this is based on the paper (Biderman et al., 2024) which shows that LoRA can mitigate forgetting; this is thus a baseline method we compare against.
@@
$\ell_{2}$: As described in (Kumar et al., 2023; Kirkpatrick et al., 2017), this is our final baseline; it advocates adding the $\ell_{2}$ distance between the old model weights and the new model weights into the loss. Note that this involves needing to store two full models in memory during training, which becomes cumbersome for large models.
Wise-FT: Taken from (Wortsman et al., 2021), this is another baseline method, based on model averaging. In particular, it advocates first doing standard fine-tuning, and then devising a weighted combination of $\alpha\times$ original model weights and $(1-\alpha)\times$ new model weights.
@@
We wish to investigate the effect of different augmentation sources on model performance. We use 1:1 mix of finetuning data and augmented samples, unless otherwise specified (we ablate on the mix of finetuning data and temperature in our experiments). For a fair comparison between different mix of datasets, we control for number of gradient steps, i.e. the size of dataset is inversely proportional to the number of epochs we use it for. We use the AdamW optimizer with a cosine learning rate schedule with a peak learning rate of 5e65e-6 and warmup steps of $3\%$ of total training steps and train for 2-4 epochs (see Supp B). We have an effective batch size of 128 (following  (Wu et al., 2025; Yuan et al., 2025)). See Supp B for additional training details. We limit our investigation to just one round of finetuning as opposed to multiple rounds with multiple different tasks. Our aim is to investigate the effect of different augmentation sources in the data-mix for continual learning necessitating making this simplifying assumption.